### v2

# GENIAL: Generative Design Space Exploration via Network Inversion for Low Power Algorithmic Logic Units

Abstract
As AI workloads proliferate, optimizing arithmetic units is becoming increasingly important for reducing the footprint of digital systems. Conventional design flows, which often rely on manual or heuristic-based optimization, are limited in their ability to thoroughly explore the vast design space. In this paper, we introduce GENIAL, a machine learning-based framework for the automatic generation and optimization of arithmetic units, with a focus on multipliers.
At the core of GENIAL is a Transformer-based surrogate model trained in two stages, involving self-supervised pretraining followed by supervised finetuning, to robustly forecast key hardware metrics such as power and area from abstracted design representations. By inverting the surrogate model, GENIAL efficiently searches for new operand encodings that directly minimize power consumption in arithmetic units for specific input data distributions. Extensive experiments on large datasets demonstrate that GENIAL is consistently more sample efficient than other methods, and converges faster towards optimized designs. This enables deployment of a high-effort logic synthesis optimization flow in the loop, improving the accuracy of the surrogate model. Notably, GENIAL automatically discovers encodings that achieve up to 18% switching activity savings within multipliers on representative AI workloads compared with the conventional two’s complement. We also demonstrate the versatility of our approach by achieving significant improvements on Finite State Machines, highlighting GENIAL’s applicability for a wide spectrum of logic functions. Together, these advances mark a significant step toward automated Quality-of-Results-optimized combinational circuit generation for digital systems.
I Introduction
This work aims to provide a modern look at how values are represented in binary format, with the ultimate goal of reducing the energy consumption of algorithmic logic units in AI accelerators and associated communication devices. An encoding is a mapping between a set of values and their corresponding binary representations. There are numerous well-recognized encodings such as the two’s complement and sign-magnitude for fixed-point numbers[arnold2025explicitsignmagnitudeencoders], the IEEE standard floating-point representations[ieee754], and the more modern Brain Float from Google[intel-bfloat16-2020, wang2019bfloat16] and HiFloat from Huawei[luo2024ascendhifloat8formatdeep]. Most of these encodings were designed to support fast and efficient binary arithmetic operations. However, they were not specifically designed to reduce internal switching activity, which is the main source of power consumption in compute-bound applications. The broadly implemented encoding for signed integer values, the Two’s Complement (TC), exhibits significant bit-level toggling for operand values close to zero and may therefore be suboptimal in terms of energy efficiency for workloads with value distributions centered around zero. The Sign-Magnitude (SM) encoding demonstrates less toggling around zero and, as a result, shows reduced switching activity for the multiplication operations[arnold2025explicitsignmagnitudeencoders]. This raises the question: are there encodings that could decrease switching activity even further than SM?
The proliferation of Large Language Models (LLM)-based chatbots and other compute-intensive generative applications is continuously increasing the global energy demand of data centers. Besides being hosted on cloud platforms, many LLMs are deployed at the edge in phones, laptops, etc. In both scenarios, quantizing weights and activations to a smaller set of integer values is a common strategy to reduce computational, communication, and memory overheads, enabling more efficient low-precision computation. Prior studies have shown that 4-bit quantization can have a negligible impact on model accuracy [NEURIPS2024_b5b93943, muller2025sinq]. Furthermore, these workloads tend to rely on numerous matrix multiplications, whose values are mostly concentrated around zero[Gongyo_2024_ACCV]. Consequently, the potential of finding better encodings for energy-efficient low-precision computations has become increasingly important.
Unlike prior approaches, such as Bayesian Optimization [grosnit2022boils] or reinforcement learning [hosny2020drills], that optimize transformation sequences or parameterized Register-Transfer Level (RTL) templates, GENIAL directly targets the operand encoding itself. This is realized through a combination of (i) self-supervised pretraining to learn robust encoding representations, (ii) legality-aware network inversion losses that preserve bijectivity, and (iii) a hierarchical EDA-in-the-loop flow with options trading off execution time and fidelity.
We propose GENIAL: a framework that enables fast and effective binary encoding discovery for optimizing circuit-level metrics such as power or area. While we mainly focus on multiplier units, we also show that our framework generalizes to other circuit types. This work brings forth the following contributions:
  • We introduce GENIAL, a machine learning-based Design Space Exploration (DSE) framework for the discovery of novel binary encodings that optimize arithmetic circuits for power efficiency.
  • We perform comprehensive ablation studies to justify each component within the framework.
  • We show that generation via Network Inversion drastically accelerates DSE, achieving superior sample efficiency compared to random or heuristic search.
  • We deploy a high-effort Logic Synthesis Optimization (LSO) flow to maximize the usefulness of the surrogate model inherent to GENIAL.
  • We conduct large-scale DSE experiments on 4-bit multipliers, where GENIAL reaches the Pareto front for Switching Activity (SwAct) and area, and discovers a new 4-bit binary encoding whose energy efficiency closely matches that of the Sign-Magnitude Extended (SME) format.
  • We release the full GENIAL framework, including training scripts, surrogate models, and EDA integration utilities, to foster reproducibility and further research. The code is available at: https://github.com/huawei-csl/GENIAL.
II Related Works
II-A Multiplier Design and Data Formats
Research on multiplier design has long centered on minimizing transistor count, reducing power consumption, and improving timing characteristics. Techniques such as Booth encoding [booth1951] and the Wallace tree [wallace1964] were among the earliest to address efficient partial product handling, and remain influential in modern designs. Over time, these ideas have evolved into higher performance variants, particularly through modifications to Booth’s original method [wen_chang_high_speed_booth]. The choice of numerical representation also has a direct impact on SwAct, which in turn affects dynamic power. TC is widely used for signed fixed-point arithmetic, but its carry chains can lead to increased bit transitions and higher dynamic power usage [SwitchinginmultipliersMsc]. The SM encoding, by comparison, mitigates this by toggling fewer bits when values are near zero [waeijen2018]. Beyond these standard formats, custom encodings tailored to arithmetic circuits have also been proposed to improve power efficiency in multipliers [costa2001power].
II-B Heuristic and Bayesian Optimization
Early efforts in design space exploration and optimization relied on exhaustive enumeration and meta‐heuristics such as simulated annealing to explore small design spaces. For instance, the binary switching algorithm, a method for optimizing mappings[64657], was applied to discover encodings that reduce bit error rates in communication systems [schreckenbach2003optimized]. More principled explorations come from Bayesian optimization (BO) with Gaussian‐process surrogates [grosnit2022boils]; later, deep neural networks were introduced to scale BO to higher dimensions [snoek2015scalable]. The BANANAS framework further improved convergence by combining BO with a path‐based neural predictor [white2021bananas].
II-C Reinforcement Learning and VAE
Deep reinforcement learning has been used to select effective sequences of logic‐synthesis transformations[hosny2020drills], and multi‐agent RL has been applied to multiplier designs[feng2024gomarl, zuo2023rl]. Variational autoencoders were also deployed for circuit interpolation[song2024circuitvae].
II-D Network Inversion
In computer vision, deep generator models synthesize inputs that maximally activate neurons [nguyen2016synthesizing], and Plug & Play Generative Networks perform iterative latent‐space optimization [nguyen2017plug]. Feature‐visualization techniques decode preferred stimuli [olah2017feature]. In protein design, inversion of AlphaFold’s structure predictor enables state-of-the-art protein sequence generation [jumper2021highly, goverde2023novo]. Collectively, these network inversion techniques demonstrate powerful generative capabilities and directly inspired GENIAL’s surrogate-based inversion approach for combinational logic.
II-E Other Hardware Optimization Methods
Another notable solution is Diffusion[NEURIPS2020_4c5bcfec], a method originally used for image generation[dhariwal2021diffusion, NEURIPS2020_4c5bcfec] and black‐box optimization[krishnamoorthy2023diffusion]. Diffusion for hardware DSE has already been adapted [ren2025diffuse] and shows promising capabilities for DSE. Finally, LLM- and agentic-based optimization has also recently attracted increasing interest. FunSearch uses Transformer agents to guide code‐generation tasks[romera-paredes_mathematical_2024], and AlphaEvolve employs LLMs with evolutionary strategies to propose design transformations[novikov2025alphaevolve]. Similarly, [yao2024rtlrewriter] optimizes RTL code. Instead of text, Circuit Transformer models directly learn on a representation of the data structures being optimized [li2024circuit, li2024logic]. Although extremely promising, LLM‐based methods for low‐level hardware synthesis remain nascent and can be costly in their execution.
III Design Space Exploration Framework
III-A Overview
GENIAL is a framework aimed at exploring the space of operand encodings to minimize specific Quality of Results (QoR) metrics of a circuit, such as area or power. As illustrated in Fig. 1, it consists of an iterative DSE process built with the following components:
1. Design Generator: converts the abstracted representation of a design to an RTL description.
2. EDA Task Launcher: a server executing Electronic Design Automation (EDA) steps such as synthesis, simulation, etc., to extract the QoR on all generated designs.
3. Quality Metric Predictor (QMP): a surrogate model trained to predict a target QoR metric for all designs out of their abstracted representations.
4. Design Recommender: a module that exploits the QMP to recommend new designs to be processed next.
At each iteration, new designs (also called prototypes) are suggested, and further processed with EDA tools. Relevant QoR metrics are extracted, and the QMP is finetuned on the entire dataset acquired so far. In this work, the term design is used to refer to all logic circuits that have the same truth table, i.e., the same binary logic function. The terms circuits or graph are used to refer to the different versions of the same design.
Figure 1: GENIAL framework overview: Generator → EDA Flows → Surrogate Model → Recommender
III-B Design Generator
III-B1 Definition of Encoding
An encoding is the mapping between a set of values and their associated binary representations. E.g., in the TC encoding, the binary representation of the value 3 is 0011 on 4 bits. A specific encoding can be specified as a dictionary $v_{i}\mapsto c_{i}$ from value to encoded representation, or as a codebook tensor (Fig. 2). There are $(2^{4})!\approx 2.1\cdot 10^{13}$ 4-bit binary encodings, which define the size of the search space for the 4-bit multiplier scenario.
Figure 2: Representation of notable encodings as 2D tensors
III-B2 Abstracted Design Representation
To interface with a machine learning pipeline, we must derive a representation of a design that can serve as input to a neural network. Since our focus is on exploring different encodings that implement the same logic function (binary multiplication), we represent each design solely by its encoding, formatted as a tensor. The columns correspond to the ordered individual bits of the binary representation, while the rows correspond to the ordered values being encoded. The TC and SME encoding representations are visible in Fig. 2.
III-B3 RTL Design Description
Generating synthesizable RTL for arbitrary operand encodings is non-trivial. To simplify this process, each design is expressed as a Look-Up-Table (LUT) circuit, allowing downstream synthesis tools to handle structural optimization. The LUT abstraction provides a technology-neutral, synthesizable representation that guarantees functional correctness for any input encoding, while outputs are consistently kept in TC format to ensure compatibility with existing digital systems. Importantly, the LUT is not the final hardware description but an intermediate form that decouples encoding legality from implementation details. Optimized RTL implementations are subsequently recovered through reliable synthesis flows such as Yosys or FLOWY. For example, a 4-bit multiplier with two inputs and one 8-bit output can be represented as:
    always @(sel) begin
        unique case(sel)
            8’b00100010 : out = 8’b00000100; // 2 x 2 = 4
        endcase
    end
As illustrated, the two 4-bit operands are concatenated into an 8-bit input pattern. All $2^{8}$ input patterns are enumerated with their corresponding 8-bit outputs.
III-C EDA Task Launcher
The task launcher performs a sequence of EDA tasks on each generated design, including synthesis, power estimation, and finally QoR metric extraction. Crucially, different metrics have different trade-offs in usefulness vs. cost: for instance, SwAct provides a close proxy to dynamic power, but requires full simulation, whereas simple complexity counts can be obtained nearly instantly. To balance accuracy and scalability, we implement three complementary flows and use them hierarchically: ESPRS for ultra-fast screening with Complexity (CPX), OPENR for mid-fidelity Number of Transistors (NT)/Power (P), and FLOWY for final SwAct/NT.
III-C1 Espresso-based Flow (ESPRS, very fast/low fidelity)
Espresso [brayton_logic_1984] is a well-established logic minimization algorithm that reduces the size of a design’s truth table. After minimization, we compute the complexity metric (CPX) by summing the counts of AND gates (in minterms) and OR gates (in maxterms). Although coarse, CPX is strongly correlated with transistor count derived by Yosys in the OPENR flow (Pearson $\rho=0.71$, see Fig. 9), making it a lightweight proxy for key QoR metrics that can be used during design space exploration.
III-C2 OpenROAD Flow (OPENR)
The OpenROAD Flow Scripts [ajayi_openroad_2019] provide a fully open-source pipeline for synthesis and place-and-route. Internally, Yosys [wolf_yosys_nodate] performs the logic synthesis, while OpenROAD [the_openroad_project_openroad_2025] and OpenSTA [the_openroad_project_opensta_2025] handle placement, routing, and timing/power analysis. The primary metrics of interest in this flow are the Number of Transistors (NT) and average Power (P), with the former being faster to compute, as it does not require simulation. NT is estimated using Yosys’ internal transistor count model. To extract P, the design is mapped to the ASAP7 predictive Process Design Kit (PDK)[CLARK2016105], followed by simulation using the open-source Verilator[verilator] simulator (v0.53). Finally, P is obtained using local SwAct data combined with PDK information. This flow strikes a balance between rapid evaluation and metric fidelity representative of hardware implementation.
III-C3 Custom Synthesis Flow (FLOWY)
We introduce a custom synthesis flow, highly inspired by the ones of [lee_late_2024, arnold_late_2025]. It consists in an initial synthesis with Yosys, succeeded by a high effort LSO using the Mockturtle (MT) [riener_scalable_2019] tool by iterative random walk, as illustrated in Fig. 3. For each generated LUT, the high effort LSO runs for 3000 steps on 12 parallel sequences of steps, called chains. At each step, a recipe is randomly selected from 30 available options and applied once if it is a decompression script, or three times if it is a compression script. At the end of the 3000 steps, the circuit with the lowest number of Majority-Inverter-Gate (MIG) gates is selected from each chain. These 12 circuits are subsequently simulated using Verilator, and the one with the lowest SwAct is selected as the final implementation for the LUT. This process is significantly slower than ESPRS or OPENR but yields highly optimized transistor-level designs that closely approximate the output of a real EDA backend.
Figure 3: Custom high-effort logic synthesis optimization flow
TABLE I: Summary of Synthesis Flows and Associated Metrics
Flow ID Method Metrics
ESPRS Espresso CPX
OPENR Open Road FlowScripts NT, P
FLOWY Yosys + MT Random Walk NT, NMIG, SwAct
III-C4 Flow Selection and Runtime Considerations
To balance accuracy and scalability, we employ a hierarchical approach using the three EDA flows introduced above:
  • ESPRS (Fast Screening): Ultra-lightweight structural proxy, allowing rapid filtering of low-quality designs based on CPX.
  • OPENR (Mid-Fidelity): Logic-synthesis pass evaluating normalized NT and coarse P, used for downselection after ESPRS.
  • FLOWY (High-Fidelity): Full LSO evaluation including SwAct, applied only to finalist candidates.
This hierarchical strategy renders end-to-end search practical by evaluating only a small subset of promising designs with the more computationally expensive flow. Table II reports typical runtime per design evaluation on our hardware (8-core CPU@3.5 GHz, no GPU acceleration for synthesis).
TABLE II: Typical runtime per design evaluation for the three EDA flows used in GENIAL.
Flow Runtime / sample #CPUs
ESPRS << 0.1 s 1
OPENR $\approx$ 1 min 1
FLOWY $\approx$ 15 min 12
In practice, ESPRS evaluates the full set of candidates generated by Network Inversion, OPENR is applied to the top fraction ranked by the surrogate model, and only a small number of finalists proceed to FLOWY for accurate power-proxy estimation.
III-C5 Switching Activity Model
For scalability, we introduce a lightweight power estimation protocol based on SwAct extraction, described below.
  1. i.
    Stimulus generation: sample an input, a 20k-long sequence of input value pairs from a normal distribution, and apply one pair of values to the multiplier’s input ports per clock cycle.
  2. ii.
    State logging: at each cycle, record the binary state of every input, output, and internal net.
  3. iii.
    Transition counting: count how many times each net state flips between consecutive cycles. Multiply each flip count by the total transistor count of that net’s fan-out cells, thereby approximating the real energy cost of an individual wire’s transitions. This is illustrated in Fig. 4.
  4. iv.
    Activity aggregation: sum all weighted costs of all nets at each clock cycle, and average over the clock cycles to derive the circuit’s average SwAct.
We validate SwAct as a proxy via strong correlations to Power and even Number of Transistors (Fig. 9). Its use in FLOWY enables accurate selection of final designs without the need to fully simulate all candidates.
Figure 4: Switching Activity model for a cell with a fanout of two
III-D Quality Metric Predictor
III-D1 Quality of Results Prediction
To recommend new designs, it is necessary to predict their QoR. Accordingly, we train a deep learning network to perform a regression on the design scores. The network takes as input the 2D tensor representation of a design’s encoding and outputs a scalar value, which corresponds to the predicted score for that encoding. Each known design is assigned a score, which is the standardized value of a target QoR metric. The encoding-score pairs serve as supervised training samples.
III-D2 Network Architecture
Fig. 5 depicts the architecture of the surrogate model, which consists in a PointNet-based[Qi_2017_CVPR] embedding followed by a stack of Transformer encoder layers[NIPS2017_3f5ee243].
Embedding: The PointNet, a sequence of one-dimensional convolution layers (C1D), converts each individual binary representation of the input encoding into a fixed-size vector, also called a token. A learned vector of the same dimension, which serves as a global representation of the design for the Transformer model and commonly called the [CLS] token, is prepended to this sequence of vectors. Finally, a learned sequence of absolute positional encodings is added element-wise to the token sequence.
Transformer and Head: The embedded token sequence is subsequently fed into a stack of Transformer encoder blocks. Each block performs non-masked multi-head self-attention and outputs a new sequence of tokens. At the output of the stack, only the vector corresponding to the [CLS] token of the final output sequence is passed through a series of linear layers, the model head, which outputs a single scalar value: the predicted score.
Figure 5: Quality Metric Predictor model architecture
III-D3 Training Pipeline
Training a network from random initialization is challenging, as it must learn both to efficiently represent each input vector in the embedding space and to predict QoR metrics from the resulting latent representation. To accelerate convergence and improve final accuracy, we adopt a two-stage training pipeline, as well as a data augmentation scheme.
Self-Supervised Pretraining: The Self-Supervised Learning (SSL) pretraining phase consists of a simple permutation prediction task. It involves predicting the positional indices of the permuted binary vectors. The TC encoding is used as the template encoding, with indices ordered from 0 to 15. For each training sample, a random permutation of the 16 binary representations and associated indices is performed. The permuted sample is passed through the Embedding and Transformer blocks. A 2-layer perceptron (MLP) receives the [CLS] token output representation and predicts the positional indices corresponding to the permutation of the binary encodings. The target is represented as a one-hot encoded sequence of length 16, and a cross entropy loss is applied. Once the SSL pretraining task is completed, the Embedding and Transformer weights are retained to initialize the model for finetuning, and the MLP head is discarded.
Finetuning: In this phase, the model is trained on the real task of predicting QoR metrics. The target score corresponds to the standardized QoR value of each known design, and the Mean Square Error (MSE) serves as the loss function. The network weights are optimized via backpropagation with the Adam optimizer [kingma2017adammethodstochasticoptimization], and the model learns to map an encoding input (represented as a 2D tensor) to its corresponding score.
Data Augmentation: We exploit some notable properties of digital designs to derive a data augmentation scheme:
  1. i.
    Swapping the positions of a logic unit’s input bits leaves its fundamental function unchanged. Similarly, permuting the columns of the encoding tensor does not change the logic function of the associated design. There are 24 possible permutations of 4 columns.
  2. ii.
    Adding inverters in front of all input bits is essentially free in terms of added logic gates. Furthermore, the SwAct depends on signal toggles, and is independent of their polarity. Hence, negating the encoding (by exchanging 0s and 1s) should likewise preserve the implementation cost.
These two transformations preserve the legality of an encoding. Thus each known design can be augmented with $2\times 24=48$ different encodings resulting in the same QoR.
III-E Design Recommendation
With the surrogate model finetuned, we leverage its predictions to guide the search toward promising designs and compare four different strategies for generating new candidates.
III-E1 Naive Approach
The Naive Approach (NA) is an intuitive baseline that involves:
  1. i.
    randomly generating a large set of candidate encodings,
  2. ii.
    predicting their score using the surrogate model, and
  3. iii.
    selecting and processing only the top designs based on predicted scores.
This process is repeated iteratively over several generations: at each round, newly generated designs are added to the existing dataset, and the surrogate model is finetuned with the updated data. While this iterative refinement is expected to progressively guide the search toward higher quality designs (QoR), in practice, the performance of this approach remains significantly worse than the Network Inversion (NI) approach described next.
III-E2 Network Inversion
The NI method employs a gradient-based optimization approach. As illustrated in Fig. 6, it proceeds by:
  1. i.
    initializing the input tensors with noise,
  2. ii.
    performing a forward pass through the network,
  3. iii.
    backpropagating gradients to the inputs, and
  4. iv.
    updating the inputs to minimize the loss.
During optimization, the network weights remain frozen; only the input tensor, representing the design encoding, is updated using the Adam [kingma2017adammethodstochasticoptimization] optimizer. The primary component of the loss is the surrogate model’s output itself, since the objective is to identify input encodings that yield low QoR values. The optimizer iteratively adjusts the input representation to reduce the loss, decreasing the surrogate prediction value. To ensure that distinct designs are generated, each input tensor is initialized with a constant value of 0.5 and perturbed with small Gaussian noise. This initialization strategy enables the generation of multiple distinct encodings while using the same fixed model parameters.
An input encoding tensor is considered legal if it contains every possible binary representation exactly once. To increase the likelihood of generating such legal encodings, we introduce two regularization terms:
  1. i.
    An attraction loss, which pulls each input vector toward all possible binary target vectors, encouraging convergence to binary values (0s and 1s).
  2. ii.
    A repulsion loss, which pushes input vectors away from one another, discouraging vectors from converging to the same target.
These regularization terms are weighted by a scalar coefficient that balances the trade-off between minimizing the QoR loss and enforcing the attraction-repulsion constraints. After 1500 optimization epochs, most of the vectors of an input converge to valid binary values, though some remain in intermediate states. A post-processing step is thus applied to legalize the encoding. Specifically, for each non-binary vector, we compute its distance to the remaining unassigned binary target vectors. The vector-target pair with the smallest distance is then matched, and both are removed from their respective candidate pools. This process is repeated iteratively until all vectors are assigned, yielding a fully legal encoding.
Figure 6: Network inversion setup
IV Experiments and Results
IV-A Experimental Setup
The experiment involves the following steps.
Model Pretraining: The model is pretrained using SSL for 200 epochs.
Initial Generation: A set of random valid designs is generated and processed to extract the QoR, forming the initial dataset.
Model Finetuning: The model is trained on the full available dataset for 200-400 epochs. The checkpoint with the lowest validation loss is selected for the recommendation.
Prototype Generation: Recommended candidate encodings are generated via NI for 1500 epochs. These 1500 NI epochs yield  99% vectors binary-converged before legalization; beyond that we saw diminishing returns.
QoR Extraction: The designs corresponding to the selected encodings are generated as LUTs. They are synthesized with one of the available flows, and their QoR are extracted.
Iteration: The newly evaluated designs are added to the dataset, and the process is repeated for another round. At each generation round, the dataset used for finetuning the network is expanded with additional designs, leading to two key effects: it improves the maximum attainable accuracy of the surrogate model, and it enriches the dataset with more diverse design knowledge. Consequently, surrogate-model-based recommendations generally improve over successive generation rounds.
Design Methodology Reference: For clarity, we briefly restate the role of each component in GENIAL:
  1. i.
    the Design Generator converts candidate encodings into synthesizable RTL (Section III-B);
  2. ii.
    the EDA Task Launcher evaluates each RTL design using one of the three flows described in Section III-C;
  3. iii.
    the Quality Metric Predictor predicts the QoR of new designs from their encoded representation (Section III-D);
  4. iv.
    the Design Recommender uses Network Inversion to propose new encodings based on surrogate gradients (Section III-E).
Figure 7: GENIAL results: OPENR with target Power (P)
IV-B OPENR Optimization for Power
In this experiment, we run GENIAL using the OPENR flow, targeting Power (P) as the QoR metric to optimize. Results are obtained by synthesizing and simulating each design with OpenROAD Flow Scripts in the synthetic ASAP7 technology [CLARK2016105]. Each design is compiled and simulated with Verilator over 20k randomly generated stimuli (i.e. input operand pair switches), with a normal distribution of mean 0 and standard deviation ($\sigma=3$). The values are clipped to the range (-8, 7) and rounded to the nearest integer. Simulation waveforms are subsequently used to perform power extraction.
Fig. 7 illustrates the evolution of the minimum and mean power (P) across generation rounds, showing that GENIAL consistently reduces power consumption over iterations. However, the LUT-based SME and TC encodings were not optimized as effectively as their hand-crafted counterparts implemented with Verilog arithmetic operators, as shown in Table III. The LUT SME design exhibits higher power than the LUT TC, which should not be the case, as highlighted by power extraction results realized on handmade designs (see Table III). Since GENIAL’s ability to discover meaningful designs strongly depends on the accuracy of the QoR metrics used to train the QMP, a higher-effort LSO in the loop is necessary.
TABLE III: OPENR Synthesis and Power Extraction Results
Metric Handmade LUT
TC SME TC SME
Area (µm2) 6.64 8.1 17.8 25.7
Power (µW) 3.25 3.14 4.52 5.34
Power vs. TC 1.00 0.97 1.00 1.18
Max Delay (ps) 1.293 1.351 1.182 1.167
Note on LUT vs. handmade RTL. As seen in Table III (OPENR), LUT-SME exhibits higher P than LUT-TC, whereas handmade operator-based RTL shows the opposite trend. We therefore treat LUT-RTL power as a screening signal only and rely on FLOWY (high-effort LSO) and SwAct for final conclusions in Sec. IV-C.
IV-C FLOWY Optimization for Switching Activity
We conduct a larger experiment with FLOWY in the loop, and the results are shown in Fig. 8. The initial dataset consists of 10k randomly generated designs, augmented with 577 additional designs uniformly sampled from a separate dataset previously generated by running GENIAL with OPENR targeting NT as the QoR metric. This augmentation increases dataset diversity and maximizes the likelihood of successful optimization. To avoid biasing the search, the initial dataset excludes both the TC or SME encodings. We apply all proposed quality-enhancement techniques, including SSL pretraining, data augmentation, and over-generation during NI, followed by a selection of the top 25% predicted encodings. For the first 10 generation rounds, 2k designs are recommended per iteration. Later, as training time became the limiting factor relative to design synthesis, we increased the number of designs per round to 4.5k. The final simulation for each best design in every chain was performed with 20k stimuli, with operand values randomly sampled from a normal distribution with $\sigma=3.0$, as above.
After generation round 13 (GR13), we observe convergence towards an encoding that exhibits significant symmetry around zero, which is different from the SME. A similar encoding was found at generation round 14 (GR14). When referring to GR13/GR14 encodings in Fig. 8, we visualize them in Fig. 2; correlations among proxies are in Fig. 9.
We find several solutions at the Pareto-front of SwAct vs NMIG. Particularly, we find solutions expected to be both smaller and to use less power than TC. Some encodings also use 12% less area than SME for only 13% more power.
We then perform a higher effort LSO (with 15k steps per chain and 120 chains per design) on selected encodings and present the results in Table IV (a). They show that the obtained encoding performs nearly as well as the SME encoding and consumes 18% less power than the TC.
Figure 8: GENIAL results: FLOWY with target SwAct
TABLE IV: FLOWY Higher Effort Synthesis Results ($\sigma=3$)
Metric Handmade LUT
TC SME TC SME GR13 GR14
Mult. (a)
NMIG 64 69 64 71 70 70
SwAct 286 231.3 321.0 235.0 263.4 272.2
vs. TC 1.00 0.73 1.00 0.73 0.82 0.85
Depth 9 12 11 9 9 9
Enc. (b)
NMIG - - 0 8 8 8
SwAct - - 0.0 34.5 38.1 39.8
Depth - - 0 3 3 3
Note that the current code implementation (available at https://github.com/huawei-csl/GENIAL) is built around Docker containers and supports multi-node, multi-CPU parallelism. This allows the search to be conducted at scale within a reasonable time frame, even when using the high-effort FLOWY synthesis (approximately one week for the full run shown in Fig. 8).
Figure 9: Correlation between various target metrics
V Validation of the Method and Discussion
V-A Metric Correlations
GENIAL is rigorously validated through a series of experiments. Since FLOWY is relatively slow (processing a single encoding takes about 15-20 min on 18 CPUs), we evaluate our methods using the faster ESPRS or OPENR flows. We observe a strong correlation between most metrics across the different flows, as shown in Fig. 9, confirming that ESPRS and OPENR flows are reliable proxies for tuning our method.
Figure 10: Comparison of Naive approach and Surrogate Model Inversion
V-B Comparison with Other Methods
V-B1 Naive versus Model Inversion
To compare NI against the NA, we evaluate both methods under equivalent compute duration, over three trials. Since NI requires several backward passes to converge on meaningful encoding prototypes, generating 10k samples takes approximately 25 minutes. Therefore, for each generation round, the NA method was run for 25 minutes, resulting in about 10M samples. From these, only the top 10k designs based on predicted scores by the QMP were retained and processed. Results of this experiment, using the OPENR flow with NT as the target QoR metric, are shown in Fig. 10. After just two generation rounds (30k processed samples), NI already discovers encodings with lower minimum NT than NA. After five rounds (60k processed samples), the NI-recommended population has a lower mean NT than the naive approach, demonstrating the superior effectiveness of NI.
Figure 11: Comparison between methods presented in section V-B2.
V-B2 Sample Efficiency
We compare our method against three others by limiting the number of processed samples across three trials. Fig. 11 shows the expected minimum value achieved as a function of the sample budget for two metrics: CPX and NT. The methods compared are:
  1. i.
    Random Search: uniformly sample legal encodings.
  2. ii.
    Binary Switching Search[64657]: starting from a random legal encoding, evaluate all swap neighbors (encodings obtained by swapping two rows in the input representation), move to the lowest‐cost neighbor (with a 25% random move chance), and restart from a new random encoding every 5000 samples processed.
  3. iii.
    Naive Approach with Surrogate Models: as described in Section III-E.
  4. iv.
    GENIAL: using 10k initial samples and 10k samples per generation.
Except at very low budgets (where the network has insufficient samples to effectively train its parameters), GENIAL consistently finds the lowest-cost encodings, demonstrating superior sample efficiency. At an equal number of processed samples, NI consistently attains lower minima and means than NA because it exploits surrogate gradients to target promising encoding manifolds directly, rather than relying on random sampling plus predictions (Fig. 11). As the dataset grows, the QMP’s accuracy improves and NI’s sample efficiency further increases.
V-C QMP Training Ablations
V-C1 SSL Pretraining
Optimizing with ESPRS, we assess the impact of SSL pretraining on recommendation quality by performing one generation round with different model checkpoints, each picked at a specific epoch during SSL. Each data point in Fig. 12 is the average of three independent runs. While higher epochs generally lead to better recommendations, no clear optimal epoch emerges.
Figure 12: Ablation study results: pretraining efforts
V-C2 Data Augmentation
Using OPENR for optimizing NT, we assess the impact of Data Augmentation (DA) during finetuning by running two experiments starting from the same dataset and pretrained models. As shown in Fig. 13, DA significantly improves recommendation quality, with both the minimum and mean NT consistently lower across all generation rounds compared to the baseline without DA.
Figure 13: Ablation study results: data-augmentation
V-D Other Considerations
V-D1 Optimizing Finite State Machines
Figure 14: Optimization results for a finite state machine
To evaluate the generalizability of our method, we applied it to Moore-type Finite State Machines (FSM), a common logic block beyond arithmetic units. We randomly generated a 16-state FSM (using 4-bit state encodings) with 3 inputs, 2 outputs, and 15 transition clauses. Starting from a random 30k-sample initial dataset and producing 5k encodings per GR with GENIAL, we observe that the minimum and mean NT are significantly lower than the initial dataset (GR0) (see Fig. 14).
Note that GENIAL is built for encoding discovery, and therefore cannot be applied to any random logic designs. GENIAL specifically targets problems with optimizable encodings.
V-D2 System-level considerations
While optimizing the operand encoding of a multiplier can reduce its power consumption, the impact must be assessed at the system level. Specifically, inputs to the multiplier, typically represented in TC, would need to be converted to a custom encoding used by the multiplier. This conversion is performed by an encoder (Enc.). Their size and cost are visible in Table IV (b). However, if the same operand can be converted once and reused multiple times, as is common in neural network accelerators, the cost of this conversion is amortized by the reuse factor.
Finally, we validated the study by implementing a DaVinci-like core [cube], which realized a General Matrix Multiplication (GEMM) operation (cube) containing $16\times 32\times 16$ multipliers. We used commercial tools for synthesis, back-annotation for glitch-aware simulation, and power extraction. By swapping out the multipliers and placing the encoders right before the matrix buffers (L0), we observed major gains:
  1. i.
    SME multipliers enable up to 16% power savings against their two’s complement counterpart on a 2D convolution operation, where both operand matrices are filled with values randomly sampled from a normal distribution with $\sigma=3.0$.
  2. ii.
    By adding the encoders’ power to the total cost, this gain is reduced to 14%.
  3. iii.
    The total area of the GEMM operator increases by 5%.
In fact, one could envision storing model weights directly in the custom encoding to further reduce system-wide latency and energy consumption.
VI Limitations
While GENIAL demonstrates that encoding-based DSE can uncover novel arithmetic designs, several limitations remain:
VI-1 Runtime of high-fidelity flows
Our framework offers multiple flows with different fidelity/cost trade-offs. Complexity (CPX) and Yosys-based transistor count (NT) can be extracted very quickly, while high-effort synthesis optimization (FLOWY) provides more realistic estimates but requires significant runtime. This limits the size of the search space that can be explored at full fidelity. In practice, we balance speed and accuracy by mixing flows depending on the experiment, but scaling to larger multipliers will require further optimization (e.g., parallelization, lightweight proxies, or adaptive flow selection).
VI-2 LUT-based design representation
All candidate multipliers are generated as LUT circuits, rather than heuristics-based or manually designed RTL implementations. This abstraction was necessary to systematically explore arbitrary input encodings while ensuring that the output remained in two’s complement, enabling compatibility with existing systems. However, LUT-based RTL can yield suboptimal intermediate results (e.g., higher switching activity compared to hand-tuned RTL) before synthesis optimization. Future work could investigate direct RTL templates or hybrid representations to close this gap.
VI-3 Scale of benchmarks
Due to the runtime and design space explosion, our experiments focus on 4-bit multipliers and randomly generated FSMs. Although there are no standard benchmarks for encoding-driven arithmetic design, further validation on larger and more diverse circuits, such as 8- or 16-bit multipliers, square-root units, and other arithmetic operators, would better demonstrate the scalability and generality of our approach.
VI-4 Exploration capability of the design recommendation
While GENIAL discovers encodings that reduce switching activity, the current exploration did not yet reach or surpass the SME format. Enhancing the search capability of the design recommendation through improved diversity control or refined objectives may help uncover superior designs. Additionally, considering alternative constraints or scenarios (e.g., multi-modal data distributions) could reveal new encodings that outperform known formats.
VI-5 Single-objective exploration guidance
The current pipeline relies on a single proxy metric to provide an optimization signal through a surrogate model. Future extensions could incorporate multiple proxies, or even an uncertainty-aware exploration strategy, enabling active sampling in regions of high predictive uncertainty. This could reveal better local minima, improve search guidance, and enhance robustness across design spaces.
Together, these limitations highlight avenues for future work, including improving runtime efficiency, exploring richer and more direct RTL representations, enhancing the diversity and multi-objective guidance of the search, and scaling evaluations to larger, more complex arithmetic benchmarks.
VII Conclusion
We introduced GENIAL, a machine-learning-driven framework for the automatic generation and optimization of combinational logic circuits. By combining a high-effort logic synthesis optimization flow with model inversion of a Transformer-based surrogate model for encoding generation, GENIAL enables sample-efficient design space exploration for optimizing key hardware metrics. Remarkably, GENIAL rediscovers the two’s-complement encoding and identifies close variants of the sign-magnitude format from scratch, without prior knowledge, while uncovering novel representations along the area–power Pareto frontier. Notably, GR13 is a 4-bit integer encoding resulting in 18% lower power consumption than the two’s complement. These findings validate surrogate-model inversion as a powerful strategy for circuit optimization. Future work will explore scaling GENIAL to larger circuits, other input value distributions, and circuits not as thoroughly investigated as multipliers, unlocking broader opportunities for logic function optimization.